# PyMieDiff: A differentiable Mie scattering library

### v2

@@ IV.2 Gradient based optimization
A central challenge in nanophotonics is designing nanostructures with desired optical properties. Design problems are typically ill-posed and need to be solved by optimization. While global optimization is useful in cases with many local minima or for non-continuous / discrete problems, this demonstraigion of local gradient based optimization provides a straightforward implementation with very fast convergence, more advanced schemes may be implemented when required.
@@
As example, we optimize the scattering efficiency $Q_{\text{sca}}$ of a dielectric core-shell particle, by modifying its core and shell size as well as the material refractive indices. We limit the radii to 1010010-100 nm and the complex refractive indexes to dielectric materials with $n_{real}$ ranging from 1 to 4.5 and $n_{imag}$ from 0.0 to 0.1. To keep the optimization within these restrictions, we optimize normalized values using a sigmoid activation inside the optimization loop.
@@
We perform a batch optimization on 100 random core-shell particles in parallel, which takes roughly $100$ ms per iteration on a typical office CPU. The spectra from the iteratively improved solutions are shown in figure 4b alongside the target spectra. The convergence curve is shown in figure 4c. We found that also higher learning rates are usually stable, allowing convergence within some 20-30 iterations. Yet as optimization is fast, lower, very stable learning rates can be used without major inconvenience.
@@ IV.3 Core-shell design: Mie informed tandem model
Our toolkit being entirely implemented in Pytorch, it can be directly implemented within any PyTorch written machine learning pipeline. Here we demonstrate how to train a deep learning neural network through a Mie-theory based loss function.
As a simple technical example, we train a design network, capable to predict core-shell geometries that fulfill given target optical properties. This so-called “tandem” model is a commonly used deep learning method for inverse design.Liu et al. (2018); Khaireh-Walieh et al. (2023) It works by regularizing the training of the ill-posed inverse problem through a forward predictor, as depicted in figure 5a. Typically, the forward model is a trained neural network that approximates a phyics solver, adding automatic differentiation capabilities. With PyMieDiff as fast, autograd-compatible Mie solver, analytical Mie theory can be used as an error-free forward model through which the loss can be backpropagated to train the inverse network.
@@
We train the Mie-informed tandem network on the same type of particles as in the optimization example, limiting the materials to constant dielectrics and training the model on spectra in the visible light range (400800400-800 nm). A dataset of 20,000 spectra, calculated from randomized particles is used for training the model. The inverse model consists of a fully connected neural network (multilayer perceptron, MLP) with three hidden layers, each containing 1024 neurons, connected by ReLU activation functions. A sigmoid function is used at the output layer to ensure positive predictions within our defined particle limits. Using a basic manually optimized training schedule, with two learning rates (first $10^{-4}$, then $10^{-5}$) and incremental increase of the batch size (from 32 up to 256 in four steps), we obtain fast convergence after a few thousand training steps (see inset in Fig. 5a). The training takes a few minutes on an 8-core office CPU.
Once the inverse model is trained, we use scattering spectra from particles that were not in the training set as test-targets and predicting the geometry parameters by the inverse network. A few examples are given in figure 5b, showing good performance of the model. Note that, as the forward model is “perfect” (exact Mie solution), the discrepancy of the solutions with respect to the traget spectra comes solely from the inverse network, for which we deliberately use a very simple dense network, chosen to have a short training time of only a few minutes.
@@ IV.4 Autodiff multi-scattering: End-to-end metasurface design
As a final example, we demonstrate how PyMieDiff can be combined with the autodiff light scattering simulation toolkit “TorchGDM”,Ponomareva et al. (2025) to perform multi-particle scattering simulations. PyMieDiff evaluates the Mie response of one or several core-shell particles, which are then used by torchGDM to generate a structure model and calculate the optical response of an ensemble of many scatterers, where all optical interactions between the particles are taken into account. Since both tools fully support PyTorch automatic differentiation, gradient optimization can be used to iteratively optimize the positions and core/shell size parameters for any target observable. This is illustrated in figure 6a.
@@
The multi-scattering capability may be useful for instance in the design of bottom-up metasurfaces or for fitting optical scattering data from dense, spherical solutions, where homogenization models fail.
@@ V Limitations and perspectives
  • PyMieDiff is currently limited to homogeneous spheres and core-shell particles. In the future we may implement multi-shell spheres as well as 2D Mie theory for infinite cylinders. The latter are more challenging to implement with pure PyTorch autodiff support, as 2D Mie theory requires cylindrical Bessel and Hankel functions, for which it is technically more difficult to implement efficient recurrences.
  • The recurrences currently require to be performed at double precision, which significantly slows down the code on consumer grade GPUs. In the future, more stable algorithms may be implemented such as logarithmic derivatives.Majic and Ru (2020); Zhang (2024)
@@
  • The current recurrences can become unstable in cases where many expansion orders are contributing to the response, for instance with very large particles or plasmonic/dielectric interfaces in larger particles. Be aware of these limitations. This could be mitigated in the future by implementing more stable algorithms, such as logarithmic derivatives recurrences.Bohren and Huffman (1998)
@@ VI Conclusions
In this work, we present PyMieDiff, an implementation of Mie scattering for core-shell particles fully built in the automatic differentiation framework PyTorch. It is available as open source software at https://github.com/UoS-Integrated-Nanophotonics-group/MieDiff. The package enables gradient-based optimisation and hybrid physics-informed deep learning models. It can be combined with other PyTorch-based toolkits, such as torchGDM, to perform end-to-end differentiable multi-particle scattering simulations. The toolkit was designed with flexibility and performance in mind, offering both, a SciPy-wrapped interface and a native PyTorch implementation with GPU support.
We demonstrate the capabilities by several examples of inverse design problems, including the iterative reconstruction of core-shell particle geometries from a target scattering spectrum, neural network training through analytical Mie calculations, and the gradient-based design of a diffractive lens made of core-shell spheres, for which PyMieDiff is combined with the multi-particle scattering toolkit “TorchGDM”.
@@ Acknowledgements.
OJ acknowledges support by EPSRC through a Ph.D. studentship. OLM acknowledges support by EPSRC through Grant No. EP/W024683/1. PRW acknowledges financial support by the French Agence Nationale de la Recherche (ANR) under grants ANR-22-CE24-0002 (project NAINOS) and ANR-23-CE09-0011 (project AIM). SDL acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 1409 published on 14/09/2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU – Project Title MINAS - CUP B53D23028420001 - Grant Assignment Decree No. 1380 adopted on 01/09/2023 by the Italian Ministry of University and Research (MUR).
@@
.1 Example code
@@
  • Second dim.: Number of particles $N_{\text{particles}}$
  • Third dim.: Number of wavenumbers $k_{0}$: $N_{\text{k0}}$
  • Fourth dim.: Number of angles or number of positions (angular scattering / fields)
Further dimensions (e.g., for field components) can be added after these four main dimensions and will be broadcasted automatically.
.3 Spherical Bessel recurrences
For comparison and testing, we provide an alternative PyTorch-based scheme to compute the spherical Bessel functions $j_{n}$ using a continued fraction calculation,Lentz (1990) Using the Riccati–Bessel form $\psi_{n}(z)=z\,j_{n}(z)$ the three-term recurrence reads
$\psi_{n+1}(z)=\frac{2n+1}{z}\,\psi_{n}(z)-\psi_{n-1}(z),$ (4)
and defining the ratio $A_{n}(z)=\psi_{n}(z)/\psi_{n-1}(z)$ gives the continued-fraction relation for stable downward ratios:
$A_{n}(z)\;=\;\frac{1}{\dfrac{2n+1}{z}-\dfrac{1}{\dfrac{2n+3}{z}-\dfrac{1}{\ddots}}}\,.$ (5)
We truncate the continued fraction at a sufficiently large $n_{\max}$,Cai (2011) compute the $A_{n}$ by backward recurrence, reconstruct $\psi_{n}$ (and hence $j_{n}=\psi_{n}/z$) up to a multiplicative constant, and fix the global scale using $j_{0}(z)=\sin z/z$. Both implementations for $j_{n}$ are similarly stable and can be selected with the argument which_jn (“ratios” or “recurrence”).
.4 Spherical Hankel functions and derivatives of spherical Bessel functions
$h^{1}_{n}(z)=j_{n}(z)-\text{i}\,y_{n}(z)\,.$ (6)
$f_{n}^{\prime}(z)$ $=f_{n-1}(z)-\frac{n+1}{z}f_{n}(z)\,$, (7)
$f_{n}^{\prime}(z)$ $=-f_{n+1}(z)+\frac{n}{z}f_{n}(z)\,$,
where $f_{n}(z)$ is any spherical bessel function.
Note: In practice, to avoid negative orders, derivatives for $n>0$ are calculated using the first relation in Eqs. (7), while the derivative for $n=0$ is obtained from the second relation:
$f_{0}^{\prime}(z)\;=\;-f_{1}(z)\,.$ (8)
@@
The vector spherical harmonics $\mathbf{N}_{lm}^{(k)}$ and $\mathbf{M}_{lm}^{(k)}$ can be used to describe the near fields of Mie scattering.Asadova et al. (2025b) For spherical symmetry, we need only the $l=1$ odd and even order harmonics, which are given by:Bohren and Huffman (1998)
$\mathbf{N}_{o1n}^{(k)}=\begin{bmatrix}\frac{1}{\rho}\sin\varphi\>n(n+1)\sin\theta\>\pi_{n}(\cos\theta)z_{n}(\rho)\\ \frac{1}{\rho}\sin\varphi\>\tau_{n}(\cos\theta)\left[\rho z_{n}(\rho)\right]^{\prime}\\ \frac{1}{\rho}\cos\varphi\>\pi_{n}(\cos\theta)\left[\rho z_{n}(\rho)\right]^{\prime}\end{bmatrix}$, (9)
$\mathbf{N}_{e1n}^{(k)}=\begin{bmatrix}\frac{1}{\rho}\cos\varphi\>n(n+1)\sin\theta\>\pi_{n}(\cos\theta)z_{n}(\rho)\\ \frac{1}{\rho}\cos\varphi\>\tau_{n}(\cos\theta)\left[\rho z_{n}(\rho)\right]^{\prime}\\ -\frac{1}{\rho}\sin\varphi\>\pi_{n}(\cos\theta)\left[\rho z_{n}(\rho)\right]^{\prime}\end{bmatrix}$, (10)
$\mathbf{M}_{o1n}^{(k)}=\begin{bmatrix}0\\ \cos\varphi\>\pi(\cos\theta)z_{n}(\rho)\\ -\sin\varphi\>\tau(\cos\theta)z_{n}(\rho)\end{bmatrix}$, (11)
$\mathbf{M}_{e1n}^{(k)}=\begin{bmatrix}0\\ -\sin\varphi\>\pi(\cos\theta)z_{n}(\rho)\\ -\cos\varphi\>\tau(\cos\theta)z_{n}(\rho)\end{bmatrix}$. (12)
Here, $m$ has become the Mie order $n$, $\rho=kr$, $\varphi$ is the azimuthal and $\theta$ the polar angle. The spherical Bessel functions of order $n$ are denoted by $z_{n}$. The kind of $z_{n}$ is indicated by the superscript (k) of $\mathbf{N}$ and $\mathbf{M}$: (1) denotes the first kind, $j_{n}(kr)$, (2) the second kind, $y_{n}(kr)$ and (3) denotes spherical Hankel functions $h_{n}^{(1)}(kr)$.
@@
The Mie coefficients for a core-shell spherical particle are given byBohren and Huffman (1998)
$a_{n}$ $\;=\;\frac{\psi_{n}(y)\left[\psi_{n}^{\prime}(m_{2}y)-A_{n}\,\chi_{n}^{\prime}(m_{2}y)\right]-m_{2}\,\psi_{n}^{\prime}(y)\left[\psi_{n}(m_{2}y)-A_{n}\,\chi_{n}(m_{2}y)\right]}{\xi_{n}(y)\left[\psi_{n}^{\prime}(m_{2}y)-A_{n}\,\chi_{n}^{\prime}(m_{2}y)\right]-m_{2}\,\xi_{n}^{\prime}(y)\left[\psi_{n}(m_{2}y)-A_{n}\,\chi_{n}(m_{2}y)\right]}$ (13)
$b_{n}$ $\;=\;\frac{m_{2}\,\psi_{n}(y)\left[\psi^{\prime}_{n}(m_{2}y)-B_{n}\,\chi^{\prime}_{n}(m_{2}y)\right]-\psi^{\prime}_{n}(y)\left[\psi_{n}(m_{2}y)-B_{n}\,\chi_{n}(m_{2}y)\right]}{m_{2}\,\xi_{n}(y)\left[\psi^{\prime}_{n}(m_{2}y)-B_{n}\,\chi^{\prime}_{n}(m_{2}y)\right]-\xi^{\prime}_{n}(y)\left[\psi_{n}(m_{2}y)-B_{n}\,\chi_{n}(m_{2}y)\right]}\,$, (14)
where $m_{1}=n_{1}/n_{\text{env}}$ and $m_{2}=n_{2}/n_{\text{env}}$ are the relative refractive indices for the core and shell, respectively and $x=ka$ and $y=kb$ are the scale factors with the wavevector in the host medium $k=k_{0}n_{\text{env}}$. The Riccati-Bessel functions are given as $\psi_{n}(z)=zj_{n}(z)$, $\xi_{n}(z)=zh_{n}^{(1)}(z)$ and $\chi_{n}(z)=zy_{n}(z)$.
$A_{n}\;=\;\frac{m_{2}\,\psi_{n}(m_{2}x)\,\psi_{n}^{\prime}(m_{1}x)-m_{1}\,\psi_{n}^{\prime}(m_{2}x)\,\psi_{n}(m_{1}x)}{m_{2}\,\chi_{n}(m_{2}x)\,\psi_{n}^{\prime}(m_{1}x)-m_{1}\,\chi_{n}^{\prime}(m_{2}x)\,\psi_{n}(m_{1}x)}\,$, (15)
and
$B_{n}\;=\;\frac{m_{2}\,\psi_{n}(m_{1}x)\,\psi_{n}^{\prime}(m_{2}x)-m_{1}\,\psi_{n}(m_{2}x)\,\psi_{n}^{\prime}(m_{1}x)}{m_{2}\,\psi_{n}(m_{1}x)\,\chi_{n}^{\prime}(m_{2}x)-m_{1}\,\psi_{n}^{\prime}(m_{1}x)\,\chi_{n}(m_{2}x)}\,.$ (16)
The Mie coefficients for internal fields can be found in literature.Bohren and Huffman (1998)
@@
The scattering, absorption and extinction efficiencies can be obtained from the Mie coefficients as,

### v3

@@ IV.2 Gradient based optimization
Design problems in nanophotonics are often non-trivial inverse problems that require numerical optimization. Here, we will try to find a core-shell particle with a response as close as possible to a hand-drawn spectrum. Since no particle likely has exactly these properties, the problem is ill-posed. Furthermore, many high-dimensional problems are naturally ill-posed, in these cases rather because several possible solutions exist.Khaireh-Walieh et al. (2023) Global optimization is useful in cases with many local minima or for non-continuous / discrete problems. Using local gradient based optimization on a large number of initial guesses provides a straightforward alternative with very fast convergence.Deng et al.
@@
In this example, we optimize the scattering efficiency $Q_{\text{sca}}$ of a dielectric core-shell particle over 400800400-800 nm to match the Gaussian centred at $600$ nm. Its core and shell size as well as the material refractive indices are varied. We limit the radii to 1010010-100 nm and the complex refractive indexes to dielectric materials with $n_{real}$ ranging from 1 to 4.5 and $n_{imag}$ from 0.0 to 0.1. To keep the optimization within these restrictions, we optimize normalized values using a sigmoid activation inside the optimization loop.
@@
We perform a batch optimization on 100 random core-shell particles in parallel, which takes roughly 100200100-200 ms per iteration on a typical office CPU. The spectra from the iteratively improved solutions are shown in figure 4b alongside the target spectra. The convergence curve is shown in figure 4c. We also found that higher learning rates are usually stable, allowing convergence within some 20-30 iterations. Yet, as optimization is fast, lower, very stable learning rates can be used without major inconvenience.
Listing 3 gives a simple code example, demonstrating how a gradient based optimization can be implemented with PyMieDiff.
@@ IV.3 Core-shell design: Mie informed tandem model
Our toolkit being entirely implemented in Pytorch, it can be directly implemented within any existing PyTorch machine learning pipeline. Here we demonstrate how to train a deep learning neural network through a Mie-theory based loss function.
As a simple technical example, we train a design network capable of predicting core-shell geometries that fulfil given target optical properties. This so called “tandem” model is a commonly used deep learning method for inverse design.Liu et al. (2018); Khaireh-Walieh et al. (2023) It works by regularizing the training of the ill-posed inverse problem through a forward predictor, as depicted in figure 5a. Typically, the forward model is a trained neural network that approximates a physics solver, adding automatic differentiation capabilities. With PyMieDiff as a fast, autograd-compatible Mie solver, analytical Mie theory can be used as an error-free forward model through which the loss can be backpropagated to train the inverse network.
@@
We train the Mie-informed tandem network on the same type of particles as in the optimization example, limiting the materials to constant dielectrics and training the model on spectra in the visible light range (400800400-800 nm). A dataset of 20,000 spectra, each consisting of 41 wavelength points, calculated from randomized particles, is used for training the model. The inverse model consists of a fully connected neural network (multilayer perceptron, MLP) with three hidden layers, each containing 1024 neurons, connected by ReLU activation functions. A sigmoid function is used at the output layer to ensure positive predictions within our defined particle limits. Using a basic manually optimized training schedule, with two learning rates (first $10^{-4}$, then $10^{-5}$) and incremental increase of the batch size (from 32 up to 256 in four steps), we obtain fast convergence after a few thousand training steps (see inset in Fig. 5a). The training takes a few minutes on an 8-core office CPU.
Once the inverse model is trained, we use scattering spectra from particles that were not in the training set as test targets and predict the geometry parameters by the inverse network. A few examples are shown in figure 5b, demonstrating the model’s performance. By comparing predicted core-shell particles to the reference particles used as design target, the results of the tandem network could now for example be used to assess “how ill posed” the studied problem of core-shell design is. The more non-unique solutions exist, the larger the discrepancy between target particles and predicted designs. Note that, as the forward model is “perfect” (exact Mie solution), the discrepancy of the solutions with respect to the target spectra comes solely from the inverse network, for which we deliberately use a very simple dense network, chosen to have a short training time of only a few minutes.
@@ IV.4 Autodiff multi-scattering: End-to-end metasurface design
As a final example, we demonstrate how PyMieDiff can be combined with the autodiff light scattering simulation toolkit “TorchGDM”,Ponomareva et al. (2025) to perform multi-particle scattering simulations. This combination enables similar functionality as offered by the T-matrix method, implemented for instance in frameworks such as CELES,Egel et al. (2017) SMARTIES,Somerville et al. (2016) or treams.Beutel et al. (2024) However, our approach enables full automatic differentiation through multi-scattering simulations.
PyMieDiff evaluates the Mie response of one or several spherical particles, which are then used by TorchGDM to generate a structure model and calculate the optical response of an ensemble of many scatterers, where all optical interactions between the particles are taken into account. Since both tools fully support PyTorch automatic differentiation, gradient optimization can be used to iteratively optimize the positions and core/shell size parameters for any target observable. This is illustrated in figure 6a.
@@
The multi-scattering capability may be useful, for instance, in the design of bottom-up metasurfaces or for fitting optical scattering data from dense, spherical solutions, where homogenization models fail.
@@ V Limitations and perspectives
  • PyMieDiff is currently limited to isotropic materials and spherical (3D) particles. In the future we may implement support for tensorial permittivities as well as 2D Mie theory for multi-shell infinite cylinders.
@@
  • While the logarithmic derivatives algorithm is very stable, there is one limitation in combination with vectorization: In a single batch of many concurrent evaluations, the Mie order required for the largest size parameter case is used for all calculations in that batch. This can in fact become unstable for small particles, causing small size parameter cases in the batch to overflow and yield NaN results (“Not a Number”) for these cases. While this is not ideal, it is failing with a non-numerical result and will directly be noted by the user. A simple split of the batch into a smaller and a larger fraction suffices to make it work.
@@ VI Conclusions
In this work, we present PyMieDiff, an implementation of Mie scattering for layered spherical particles, fully built in the automatic differentiation framework PyTorch. It is available as open source software at https://github.com/UoS-Integrated-Nanophotonics-group/MieDiff. The package enables gradient-based optimisation and hybrid physics-informed deep learning models. It can be combined with other PyTorch-based toolkits, such as torchGDM, to perform end-to-end differentiable multi-particle scattering simulations. The toolkit was designed with flexibility and performance in mind, offering a native PyTorch implementation of the entire Mie algorithm, with full GPU support.
We demonstrate the capabilities by several examples of inverse design problems, including the iterative reconstruction of core-shell particle geometries from a target scattering spectrum, neural network training through analytical Mie calculations, and the gradient-based design of a diffractive lens made of several core-shell spheres, for which PyMieDiff is combined with the multi-particle scattering toolkit “TorchGDM”.
@@ Acknowledgements.
OJ acknowledges support by EPSRC through a PhD studentship. OLM acknowledges support by EPSRC through Grant No. EP/W024683/1. PRW acknowledges financial support by the French Agence Nationale de la Recherche (ANR) under grants ANR-22-CE24-0002 (project NAINOS) and ANR-23-CE09-0011 (project AIM). SDL acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 1409 published on 14/09/2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU – Project Title MINAS - CUP B53D23028420001 - Grant Assignment Decree No. 1380 adopted on 01/09/2023 by the Italian Ministry of University and Research (MUR).
@@
.1 Stability on very large spheres
Figure A1: Demonstration of stability on large spheres compared to scattnlay.Peña and Pal (2009) (a) Scattering of a soot-coated water sphere in vacuum for increasing size parameter. Used refractive indixes are $n_{\text{water}}=1.33$ and $n_{\text{soot}}=1.59+0.66\,\text{i}$. The soot volume fraction is $0.01$. (b) Angular scattering from a layered sphere with layer radii $[135,2365,2395,15000]$ nm and layer refractive indices $2.1+0.15\text{i},1.75,0.45+5.06\text{i},3.62]$ at $\lambda_{0}=1100$ nm.
To assess the stability of PyMieDiff on very large problems, we reproduce the test performed by Peña and Pal,Peña and Pal (2009) who simulate a soot shell around a large water sphere for increasing size parameters. The result is shown in Figure A1a.
In Figure A1b we show angular scattering for a 4-layer particle of 3030\,µm outer diameter, consisting mainly of dielectric material but containing also absorbing shells. The size parameter of this problem is 310.
.2 Example code
@@
  • Second dim.: Layer number $l$: l=1,,Ll=1,...,L
  • Third dim.: Number of particles $N_{\text{particles}}$
  • Fourth dim.: Number of wavenumbers $k_{0}$: $N_{\text{k0}}$
  • Fith dim.: Number of angles or number of positions (angular scattering / fields)
Further dimensions (e.g., for field components) can be added after these four main dimensions and will be broadcast automatically.
.4 Riccati–Bessel recurrences
To ensure numerical stability, we implement the logarithmic derivatives of the Riccati–Bessel functions, $D_{n}^{(1)}(z)$ and $D_{n}^{(3)}(z)$, rather than their standard derivative, $\psi_{n}^{\prime}(z)$ and $\xi_{n}^{\prime}(z)$ Yang (2003). The logarithmic derivative of the first kind, $D_{n}^{(1)}(z)=\psi_{n}^{\prime}(z)/\psi_{n}(z)$, is evaluated using the downward recurrence Wiscombe (1980):
$D_{n-1}^{(1)}(z)=\frac{n}{z}-\frac{1}{D_{n}^{(1)}+\frac{n}{z}},\quad n=N_{max},\dots,1,$ (3)
with the initial condition:
$D_{N_{max}}^{(1)}(z)=0+i0.$ (4)
Conversely, the logarithmic derivative of the third kind, $D_{n}^{(3)}(z)=\xi_{n}^{\prime}(z)/\xi_{n}(z)$, stability across all values of z is achieved by utilizing an upward recurrence relation Mackowski et al. (1990):
$D_{n}^{(3)}(z)=D_{n}^{(1)}(z)+\frac{i}{\psi_{n}(z)\xi(x)},\quad n=1,\dots,N_{max},$ (5)
This requires an auxiliary upward recurrence to compute the product $\psi_{n}(z)\xi_{n}(z)$:
$\psi_{n}(z)\xi_{n}(z)=\psi_{n-1}(z)\xi_{n-1}(z)\left[\frac{n}{z}-D_{n-1}^{(1)}(z)\right]\left[\frac{n}{z}-D_{n-1}^{(3)}(z)\right],$ (6)
which is initialized as follows:
$D_{0}^{(3)}(z)=i,$ (7)
$\psi_{0}(z)\xi_{0}(z)=\frac{1}{2}[1-(\cos 2a+i\sin 2a)\exp(-2b)].$ (8)
The plain Riccati–Bessel functions are found using the recurrence relations Mackowski et al. (1990); Wiscombe (1980):
$\psi_{n}\left(z\right)=\psi_{n-1}\left(z\right)\left[\frac{n}{z}-D_{n-1}^{(1)}\left(z\right)\right],\quad n=1,\ldots,N_{\max},$ (9)
$\xi_{n}\left(x_{L}\right)=\xi_{n-1}\left(z\right)\left[\frac{n}{z}-D_{n-1}^{(3)}\left(z\right)\right],\quad n=1,\ldots,N_{\max},$ (10)
starting from,
$\psi_{0}\left(z\right)=\sin\left(z\right),$ $\xi_{0}\left(z\right)=\sin\left(z\right)-i\cos\left(z\right),$ (11)
respectfully. The ideal $N_{max}$ and $N_{stop}$ are found programmatically based of the work of Wiscombe.Wiscombe (1980)
@@
The vector spherical harmonics $\mathbf{N}_{lm}^{(k)}$ and $\mathbf{M}_{lm}^{(k)}$ can be used to describe the near fields of Mie scattering.Asadova et al. (2025b) For spherical symmetry, only the $l=1$ odd and even order harmonics are required, which can be written using the Riccati–Bessel functions and their logarithmic derivatives:Ladutenko et al. (2017)
$\mathbf{N}_{o1n}^{(j)}=\begin{bmatrix}\sin\phi\;n(n+1)\sin\theta\;\pi_{n}(\cos\theta)\frac{z_{n}(\rho)}{\rho^{2}}\\ \sin\phi\;\tau_{n}(\cos\theta)\frac{D_{n}^{(j)}(\rho)z_{n}(\rho)}{\rho}\\ \cos\phi\;\pi_{n}(\cos\theta)\frac{D_{n}^{(j)}(\rho)z_{n}(\rho)}{\rho}\end{bmatrix}$, (12)
$\mathbf{N}_{e1n}^{(j)}=\begin{bmatrix}\cos\phi\;n(n+1)\sin\theta\;\pi_{n}(\cos\theta)\frac{z_{n}(\rho)}{\rho^{2}}\\ \cos\phi\;\tau_{n}(\cos\theta)\frac{D_{n}^{(j)}(\rho)z_{n}(\rho)}{\rho}\\ -\sin\phi\;\pi_{n}(\cos\theta)\frac{D_{n}^{(j)}(\rho)z_{n}(\rho)}{\rho}\end{bmatrix}$, (13)
$\mathbf{M}_{o1n}^{(k)}=\begin{bmatrix}0\\ \cos\phi\;\pi_{n}(\cos\theta)\frac{z_{n}(\rho)}{\rho}\\ -\sin\phi\;\tau_{n}(\cos\theta)\frac{z_{n}(\rho)}{\rho}\end{bmatrix}$, (14)
$\mathbf{M}_{e1n}^{(k)}=\begin{bmatrix}0\\ -\sin\phi\;\pi_{n}(\cos\theta)\frac{z_{n}(\rho)}{\rho}\\ -\cos\phi\;\tau_{n}(\cos\theta)\frac{z_{n}(\rho)}{\rho}\end{bmatrix}$, (15)
Here, $m$ has become the Mie order $n$, $\phi$ is the azimuthal and $\theta$ the polar angle. The Riccati–Bessel functions of order $n$ are denoted by $z_{n}$. The kind of $z_{n}$ is indicated by the superscript (k) of $\mathbf{N}$ and $\mathbf{M}$: (1) denotes $\psi_{n}$ while (3) denotes $\xi_{n}$, and their corresponding logarithmic derivatives, $D_{n}^{(1)}$ and $D_{n}^{(3)}$. These functions are given by $\psi_{n}(\rho)=\rho j_{n}(\rho)$ and $\xi_{n}(\rho)=\rho h_{n}^{(1)}(\rho)$.
@@
The formula of the scattering coefficients for a spherical particle with $L$ layers first developed by YangYang (2003) and later modernized by Peña and Pal Peña and Pal (2009) are given by:
$a_{n}=a_{n}^{L+1}=\frac{\left[H_{n}^{a}\left(m_{L}x_{L}\right)/m_{L}+n/x_{L}\right]\psi_{n}\left(x_{L}\right)-\psi_{n-1}\left(x_{L}\right)}{\left[H_{n}^{a}\left(m_{L}x_{L}\right)/m_{L}+n/x_{L}\right]\xi_{n}\left(x_{L}\right)-\xi_{n-1}\left(x_{L}\right)},$ (16)
$b_{n}=b_{n}^{L+1}=\frac{\left[m_{L}H_{n}^{b}\left(m_{L}x_{L}\right)+n/x_{L}\right]\psi_{n}\left(x_{L}\right)-\psi_{n-1}\left(x_{L}\right)}{\left[m_{L}H_{n}^{b}\left(m_{L}x_{L}\right)+n/x_{L}\right]\xi_{n}\left(x_{L}\right)-\xi_{n-1}\left(x_{L}\right)},$ (17)
where $m_{l}$ is the relative refractive index and $x_{l}$ is the size parameter of each layer, $l$ with l=1,,Ll=1,...,L. The size parameter for each layer is given by $x_{l}=kr_{l}$ with the wavevector in the host medium $k=k_{0}n_{\text{env}}$ and the radii of layer $r_{l}$. $\psi_{n-1}(x_{L})$ and $\xi_{n-1}(x_{L})$ are found using the reccurrence relation outlined in Section .4, while the composite functions $H^{a}_{n}$ and $H^{b}_{n}$ are given by:
$H_{n}^{a}\left(m_{1}x_{1}\right)=D_{n}^{(1)}\left(m_{1}x_{1}\right),$ (18)
$H_{n}^{a}\left(m_{l}x_{l}\right)=\frac{G_{2}D_{n}^{(1)}\left(m_{l}x_{l}\right)-Q_{n}^{(l)}G_{1}D_{n}^{(3)}\left(m_{l}x_{l}\right)}{G_{2}-Q_{n}^{(l)}G_{1}},\quad l=2,\ldots,L,$ (19)
$H_{n}^{b}\left(m_{1}x_{1}\right)=D_{n}^{(1)}\left(m_{1}x_{1}\right),$ (20)
$H_{n}^{b}\left(m_{l}x_{l}\right)=\frac{\widetilde{G}_{2}D_{n}^{(1)}\left(m_{l}x_{l}\right)-Q_{n}^{(l)}\widetilde{G}_{1}D_{n}^{(3)}\left(m_{l}x_{l}\right)}{\widetilde{G}_{2}-Q_{n}^{(l)}\widetilde{G}_{1}},\quad l=2,\ldots,L,$ (21)
where,
$G_{1}=m_{l}H_{n}^{a}\left(m_{l-1}x_{l-1}\right)-m_{l-1}D_{n}^{(1)}\left(m_{l}x_{l-1}\right),$ (22)
$G_{2}=m_{l}H_{n}^{a}\left(m_{l-1}x_{l-1}\right)-m_{l-1}D_{n}^{(3)}\left(m_{l}x_{l-1}\right),$ (23)
$\widetilde{G}_{1}=m_{l-1}H_{n}^{b}\left(m_{l-1}x_{l-1}\right)-m_{l}D_{n}^{(1)}\left(m_{l}x_{l-1}\right),$ (24)
$\widetilde{G}_{2}=m_{l-1}H_{n}^{b}\left(m_{l-1}x_{l-1}\right)-m_{l}D_{n}^{(3)}\left(m_{l}x_{l-1}\right).$ (25)
These are found recursively, details on $D_{n}^{(1)}$, $D_{n}^{(3)}$ are oulined in Section .4, while the ratio:
$Q_{n}^{(l)}=\frac{\psi_{n}\left(m_{l}x_{l-1}\right)}{\xi_{n}\left(m_{l}x_{l-1}\right)}/\frac{\psi_{n}\left(m_{l}x_{l}\right)}{\xi_{n}\left(m_{l}x_{l}\right)},$ (26)
can be found via the upwards recccurence relation Yang (2003):
$Q_{n}^{(l)}=Q_{n-1}^{(l)}\left(\frac{x_{l-1}}{x_{l}}\right)^{2}\frac{\left[z_{2}D_{n}^{(1)}\left(z_{2}\right)+n\right]}{\left[z_{1}D_{n}^{(1)}\left(z_{1}\right)+n\right]}\frac{\left[n-z_{2}D_{n-1}^{(3)}\left(z_{2}\right)\right]}{\left[n-z_{1}D_{n-1}^{(3)}\left(z_{1}\right)\right]},$ (27)
for $n=1,\ldots,N_{\max}$, with the initial condition:
$Q_{0}^{(l)}=\frac{\exp\left(-i2a_{1}\right)-\exp\left(-2b_{1}\right)}{\exp\left(-i2a_{2}\right)-\exp\left(-2b_{2}\right)}\times\exp\left(-2\left[b_{2}-b_{1}\right]\right),$ (28)
where $z_{1}=m_{l}x_{l-1}=a_{1}+ib_{1}$ and $z_{2}=m_{l}x_{l}=a_{2}+ib_{2}$.
@@
The scattering, absorption and extinction efficiencies can be obtained from the scattering coefficients as,